# Embodying Control in Soft Multistable Grippers from Morphofunctional Co-design

### v2

@@ 3 Conclusion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics, discretizing their continuous response using multistable structures. We propose a robotic architecture, The Dome Phalanx Finger, that leverages mechanical multistability to discretize the robots’ configuration space into desired kinematics, which serves as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses providing an alternative open-loop control strategy for realizing various robotic tasks normally requiring closed-loop approaches guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to accurately predict and program the DPF’s behavior. Through this approach, we established a robust framework for modeling the finger’s static and dynamic responses, with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and recursive feature elimination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. Our results show how these characteristics can be designed to allow for the modulation of grasping force and unit reset without the need for closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. More importantly, this approach is not only beneficial for systems controlled in open loop. As the robot behavior is transformed into a finite, tractable set of positions or set points, the implementation of closed-loop control strategies is simplified by the tracking of a finite number of states instead of the infinite number inherent to continuous systems. Combining our morphological control strategy with conventional sensors can create a streamlined closed-loop control, where sensors can be utilized to verify if the robot is in a specific state, thus reducing the need for constant feedback from the system. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
@@ 4 Experimental Section
Unit cell and DPF Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [52] and Ninjatek Cheetha TPU [53]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 6 for assembly).
Stiffness Measurement tests– The actuators were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Figure S16a and Movie 5). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
@@ A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential dome units (Figure S1d). The finger derives its multistable behavior from domed-shaped shell elements (Figure S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{\text{lim}}$), air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$), spacing between adjacent cells ($\text{U}^{i}_{\text{sep}}$), unit cell length ($\text{U}^{i}_{\text{L}}$), chamber thickness ($t_{ch}$) and dome thickness. Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Figure S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed. The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Figure S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{ch}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Figure S2b) and the final tip position of the system (Figure S2c).
@@
Consequently, various dynamic behaviors can be achieved by adjusting the geometrical parameters, thereby enhancing tunability and control. Specifically, the dome height ($H$) and thickness ($t$), along with the limiting layer thickness ($t_{\text{lim}}$), influence the final stable position or set point. Meanwhile, the chamber thickness ($t_{ch}$) and the air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$) govern the overshoot required to reach a target position (see Figure S3). This tunability enables the gripper’s response to be tailored by simply modifying the DPF’s geometric characteristics.
@@ A.3.1 Spring Lattice array
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising nonlinear [49], linear, and torsional springs (Figure S5a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use nonlinear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the nonlinear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the nonlinear units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Figure S5a).We define $k_{l}=\frac{\sqrt{3}}{2}E\frac{t_{\text{lim}}UC}{U_{L}+t_{\text{ch}}+U_{\text{sep}}}$ and $k_{t}=\frac{2}{\sqrt{3}}\frac{E(t_{lim}+t)}{12(1-\nu^{2})}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Figure S2a.
@@ A.3.3 Parameter Tuning
$\begin{split}k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{E}{R^{2}}\Bigl{(}-\frac{1.97Rt^{5}}{H^{3}}+\frac{7.4H^{2}t^{3}}{R^{2}}+\frac{3.5Rt^{4}}{H^{2}}+\frac{0.37H^{2}t^{2}}{R}+\frac{42.2t^{5}}{H^{2}}\\ &-\frac{35.8Ht^{4}}{R^{2}}+\frac{71.8t^{5}}{HR}-\frac{3.4Rt^{3}}{H}-\frac{67.7t^{4}}{H}+4.2Rt^{2}+11.1t^{3}\Bigr{)}\end{split}$ (S.16)
$\begin{split}\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{0.29H^{3}}{R^{3}}-\frac{0.44t^{3}}{H^{3}}-\frac{2.2H^{2}t}{R^{3}}-\frac{0.73H^{2}}{R^{2}}-\frac{6.9t^{3}}{H^{2}R}+\frac{1.17t^{2}}{H^{2}}-\frac{3.0t^{3}}{HR^{2}}+\\ &\frac{7.05Ht}{R^{2}}+\frac{13.38t^{2}}{HR}+\frac{0.44H}{R}+\frac{13.67t^{3}}{R^{3}}-\frac{15.0t^{2}}{R^{2}}-\frac{4.7t}{R}\end{split}$ (S.17)
@@
A.4 Static and Dynamic Response
@@
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the viscoelastic response of the material. By incorporating these elements and using symbolic differentiation, a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{x}_{i}+F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)=F_{ext}(x_{i},t)$ (S.19)
where $[M]$ is the mass matrix of the system, $F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})$ is the internal damping, which can be modeled as a dashpot in parallel to each spring (see Figure 3a) and $F_{d}^{iso}(\dot{x}_{i})$ is the isotropic drag, which acts as an external force on each node proportional to velocity ($\dot{x}_{i}$[55]. $F_{ext}(x_{i,j},t)$ is the external actuation pressure that can be calculated as $F_{ext}(x_{ij},t)=A(x_{ij},)P(t)$. Where $A(x_{ij},)$ is the area of the face where the pressure is applied. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{x}_{i}$, we can rewrite the system as
@@
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. We explore the dynamic behavior of a five-segment DPF by applying the internal pressure as shown in Figure S7c, where we obtain the final stable state as shown and the dynamic response shown in Figure S7d. As expected, the model can capture the dynamic response of a bistable, metastable, and monostable DPF, where snap-through and snap-back phenomena are observed.
Furthermore, we analyze the temporal response of the metastable units to assess how geometrical parameters and material damping influence the resetting time. Our results show that adjusting the dome unit’s thickness modifies the resetting time (see Figure S9b). However, an even more pronounced effect is observed when varying the loading duration (see Figure S8c). This suggests that after manufacturing the DPF, additional tunability can be achieved by controlling the applied pressure (see Movie 7).
@@
The resetting time response is validated against experimental data to assess its predictive accuracy. As shown in Figure S9c, the model effectively captures the time scales of a fully metastable DPF. However, due to the inherent randomness in the resetting process when all units are metastable, the tip displacement is not precisely predicted. Nevertheless, this tool enables the design of both the static and dynamic responses by programming the resetting time scales of the metastable domes (see Movie 2 for experimental comparison).

### v3

@@ 3 Conclusion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics enabling the embodying of their control by discretizing their continuous response using multistable structures. We propose two robotic architectures, The Dome Phalanx Gripper and Dome Phalanx Walker, that leverage mechanical instabilities to discretize the robots’ configuration space into desired kinematics, which serve as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses, providing an alternative open-loop control strategy for realizing various robotic tasks typically requiring closed-loop approaches guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to accurately predict and program the DPF’s behavior and units. This approach established a robust framework for modeling the finger’s static and dynamic responses with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and recursive feature elimination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality, yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. More importantly, we show that we can perform simple classification tasks by combining the robot morphology and the simple feedback. We demonstrate that multistability enables for damage-tolerance in inflatable robots and shape and dynamics are less reliant on careful pressurization. Furthermore, the versatility of our approach is illustrated by enabling the controlled locomotion and steering of a walker by open-loop simple pressure modulation. As the robot behavior is transformed into a finite, tractable set of positions or set points, the implementation of closed-loop control strategies is simplified by tracking a finite number of states instead of the infinite number inherent to continuous systems. Combining our morphological control strategy with conventional sensors can create a streamlined closed-loop control, where sensors can be utilized to verify if the robot is in a specific state, thus reducing the need for constant feedback from the system. Our results show how these characteristics can be designed to allow for dynamic reconfiguration, grasping force modulation, and embodied logic in the robot morphology without needing closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
@@ 4 Experimental Section
Unit cell and DPF Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [52] and Ninjatek Cheetha TPU [53]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 5 for assembly).
Stiffness Measurement tests– The fingers were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Figure S15a). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
@@ A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential dome units (Figure S1d). The finger derives its multistable behavior from domed-shaped shell elements (Figure S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{\text{lim}}$), air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$), spacing between adjacent cells ($\text{U}^{i}_{\text{sep}}$), unit cell length ($\text{U}^{i}_{\text{L}}$), chamber thickness ($t_{ch}$) and dome thickness. Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Figure S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed.
@@
The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Figure S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{ch}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Figure S2b) and the final tip position of the system (Figure S2c). Consequently, various dynamic behaviors can be achieved by adjusting the geometrical parameters, thereby enhancing tunability and control. Specifically, the dome height ($H$) and thickness ($t$), along with the limiting layer thickness ($t_{\text{lim}}$), influence the final stable position or set point. Meanwhile, the chamber thickness ($t_{ch}$) and the air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$) govern the overshoot required to reach a target position (see Figure S3). This tunability enables the gripper’s response to be tailored by simply modifying the DPF’s geometric characteristics.
@@ A.3.1 Spring Lattice array
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising nonlinear [49], linear, and torsional springs (Figure S5a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use nonlinear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the nonlinear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the nonlinear units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Figure S5a).We define $k_{l}=\frac{Et_{\text{lim}}UC}{U_{L}}$ and $k_{t}=\frac{Et^{3}_{lim}}{12(1-\nu^{2})}\frac{UC}{U_{\text{sep}}}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Figure S2a.
@@ A.3.3 Parameter Tuning
$k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)$ $=\frac{E}{R^{2}}\Biggl{(}-\frac{1.97Rt^{5}}{H^{3}}+\frac{7.4H^{2}t^{3}}{R^{2}}+\frac{3.5Rt^{4}}{H^{2}}+\frac{0.37H^{2}t^{2}}{R}+\frac{42.2t^{5}}{H^{2}}$ (S.15)
$\quad-\frac{35.8Ht^{4}}{R^{2}}+\frac{71.8t^{5}}{HR}-\frac{3.4Rt^{3}}{H}-\frac{67.7t^{4}}{H}+4.2Rt^{2}+11.1t^{3}\Biggr{)}$
$\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)$ $=\frac{0.4H^{3}}{R^{3}}-\frac{0.6t^{3}}{H^{3}}-\frac{2.9H^{2}t}{R^{3}}-\frac{0.9H^{2}}{R^{2}}-\frac{9.0t^{3}}{H^{2}R}+\frac{1.5t^{2}}{H^{2}}$ (S.16)
$\quad-\frac{3.9t^{3}}{HR^{2}}+\frac{9.2Ht}{R^{2}}+\frac{17.4t^{2}}{HR}+\frac{0.6H}{R}$
$\quad+\frac{17.8t^{3}}{R^{3}}-\frac{19.5t^{2}}{R^{2}}-\frac{6.1t}{R}.$
@@
A.3.4 Derivation of Jacobian terms
Here we detail the derivation for the energy terms of the linear. non-linear and rotational springs.
Linear Springs: The force on a node $i$ at position $x_{i}$ interacting with and adjacent node $j$ at a position $x_{j}$ is given by:
$F_{L}(x_{i},x_{j})=\nabla_{x_{i}}E_{L}=\frac{1}{2}k_{l}\nabla_{x_{i}}\left(||x_{ij}||-s_{ij}\right)^{2}=k_{l}\left(\frac{s_{ij}}{||x_{ij}||}-1\right)x_{ij}$ (S.18)
where $x_{ij}=x_{i}-x_{j}$ and $s_{ij}$ is the rest length of the spring.
Nonlinear Springs: The term term of adjacent nodes $i$ and $j$ connected by a bistable spring can be written as:
$E_{NL}=\frac{1}{2}k_{b}\bar{x}^{2}\left(1+(1-\alpha)\left(\frac{\bar{x}^{2}}{d^{2}}-2\frac{\bar{x}}{d}\right)\right)$ (S.19)
where $\bar{x}=||x_{ij}||-s_{ij}$. Given this, the non-linear spring forces can be written as:
$F_{NL}(x_{i},x_{j})=\nabla_{x_{i}}E_{NL}=k_{b}\bar{x}\left(1+(1-\alpha)\left(2\frac{\bar{x}^{2}}{d^{2}}-3\frac{\bar{x}}{d}\right)\right)\frac{x_{ij}}{||x_{ij}||}$ (S.20)
Rotational Springs: The force on a node $i$ due to bending at the limiting layer is given by
$F_{T}=\nabla_{x_{i}}E_{T}(\vartheta)=k_{\vartheta}\left(\vartheta-\vartheta_{0}\right)\nabla_{x_{i}}\vartheta$ (S.21)
This force is an interaction between three nodes, thus the gradient $\nabla_{x_{i}}$ is computed with respect to $x_{i}\in{x_{0},x_{1},x_{2}}$. The relevant vectors can be written as
$e_{0}=x_{1}-x_{0}$ (S.22)
$e_{1}=x_{2}-x_{0}$ (S.23)
with $\hat{e}_{i}=e_{i}/||e_{i}||$ denotating the normal vector. We can calculate the gradient of $\vartheta$ as
$\nabla_{x_{i}}\vartheta=\nabla_{x_{i}}\left(\cos^{-1}\left(\hat{e}_{0}\cdot\hat{e}_{1}\right)\right)=-\frac{\nabla_{x_{i}}\left(\hat{e}_{0}\cdot\hat{e}_{1}\right)}{\sin\vartheta}$ (S.24)
The gradients with respect to each component of $x_{i}$ may be expressed as
$\nabla_{x_{0}}\left(\hat{e}_{0}\cdot\hat{e}_{1}\right)=-\frac{e_{0}+e_{1}}{||e_{0}||||e_{1}||}+\left(e_{0}\cdot e_{1}\right)\left(\frac{e_{0}}{||e_{0}||^{3}||e_{1}||}+\frac{e_{1}}{||e_{0}||||e_{1}||^{3}}\right)$ (S.25)
$\nabla_{x_{1}}\left(\hat{e}_{0}\cdot\hat{e}_{1}\right)=\frac{1}{||e_{0}||||e_{1}||}\left(e_{1}-\frac{e_{0}\cdot e_{1}}{||e_{0}||^{2}}e_{0}\right)$ (S.26)
$\nabla_{x_{2}}\left(\hat{e}_{0}\cdot\hat{e}_{1}\right)=\frac{1}{||e_{0}||||e_{1}||}\left(e_{0}-\frac{e_{0}\cdot e_{1}}{||e_{1}||^{2}}e_{1}\right)$ (S.27)
Using this expression, we can calculate $\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)=F_{L}+F_{NL}+F_{T}$.
A.4 DPF Static and Dynamic Response
@@
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the viscoelastic response of the material. By incorporating these elements and determining its jacobian terms (see section A.3.4), a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{x}_{i}+F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)=F_{ext}(x_{ij},t)$ (S.28)
where $[M]$ is the mass matrix of the system, $F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})$ is the internal damping, which can be modeled as a dashpot in parallel to each spring (see Figure 3a) and $F_{d}^{iso}(\dot{x}_{i})$ is the isotropic drag, which acts as an external force on each node proportional to velocity ($\dot{x}_{i}$[55]. $F_{ext}(x_{i,j},t)$ is the external actuation pressure that can be calculated as $F_{ext}(x_{ij},t)=A(x_{ij})P(t)$. Where $A(x_{ij},)$ is the area of the face where the pressure is applied. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{x}_{i}$, we can rewrite the system as
@@
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. We explore the dynamic behavior of a five-segment DPF by applying the internal pressure as shown in Figure S7c, where we obtain the final stable state as shown and the dynamic response shown in Figure S7d. As expected, the model can capture the dynamic response of a bistable, metastable, and monostable DPF, where snap-through and snap-back phenomena are observed. Furthermore, we analyze the temporal response of the metastable units to assess how geometrical parameters and material damping influence the resetting time. Our results show that adjusting the dome unit’s thickness modifies the resetting time (see Figure S9b). However, an even more pronounced effect is observed when varying the loading duration (see Figure S8c). This suggests that after manufacturing the DPF, additional tunability can be achieved by controlling the applied pressure (see Movie 6). The resetting time response is validated against experimental data to assess its predictive accuracy. As shown in Figure S9c, the model effectively captures the time scales of a fully metastable DPF. However, due to the inherent randomness in the resetting process when all units are metastable, the tip displacement is not precisely predicted. Nevertheless, this tool enables the design of both the static and dynamic responses by programming the resetting time scales of the metastable domes (see Movie 2 for experimental comparison).